OpenStudy (anonymous):

Are 20, 25 and 30 an example of a Pythagorean triplet? Yes or No, Explain your answer and show work.

OpenStudy (xapproachesinfinity):

Pythagorean triples are the numbers that satisfy \(\large a^2+b^2=c^2\) which the Pythgoras theorem the numbers a,b,c are called Pythagorean triple compare this to what you have

OpenStudy (xapproachesinfinity):

to do so you need to find what is \(\large 20^2+25^2=?\)

OpenStudy (xapproachesinfinity):

then compare that to \(\large 30^2\)